# On the Anti-Ramsey Number Under Edge Deletion

### v1

@@ Abstract
According to a study by Erdős et al. in 1975, the anti-Ramsey number of a graph $G$, denoted as $AR(n,G)$, is defined as the maximum number of colors that can be used in an edge-coloring of the complete graph $K_{n}$ without creating a rainbow copy of $G$. In this paper, we investigate the anti-Ramsey number under edge deletion and demonstrate that both decreasing and unchanging are possible outcomes. For three non-negative integers $k$, $t$, and $n$, let $G=kP_{4}\cup tP_{2}$. Let $E^{\prime}$ be a subset of the edge set $E(G)$ such that every endpoint of these edges has a degree of two in $G$. We prove that if one of the conditions (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; (iii) $k=1$, $t\geq 1$, and $n\geq 2t+4$, occurs then the behavior of the anti-Ramsey number remains consistent when the edges in $E^{\prime}$ are removed from $G$, i.e., $AR(n,G)=AR(n,G-E^{\prime})$. However, this is not the case when $k\geq 2$, $t=0$, and $n=4k$. As a result, we calculate $AR(kP_{4}\cup tP_{2})$ for the cases: (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; (iii) $k=1$, $t\geq 0$, and $n\geq 2t+4$; (iv) $k\geq 1$, $t=0$, and $n=4k$.
@@ 1 Introduction
For a positive integer $n$ and a graph $H$, the anti-Ramsey number $AR(n,H)$ is defined as the maximum number of colors in an edge-coloring of the complete graph $K_{n}$ that does not contain a rainbow copy of $H$. This concept was introduced by Erdős et al. in 1975 [3], where they established its connection to Turán numbers and conjectured values for the anti-Ramsey numbers of paths and cycles. Subsequent research has extensively explored anti-Ramsey numbers for graphs consisting of multiple small, disjoint components. A foundational result is the identification of the anti-Ramsey number for matchings (disjoint edges) in complete graphs [2, 5, 8, 10, 19]. More recently, the focus has shifted to more complex structures, particularly the vertex-disjoint union of small connected graphs and matchings. Our work contributes to this ongoing inquiry by studying linear forests composed of short odd paths. To clarify the findings related to our study so far, we outline them below. Simonovits and Sós [18] proved that if $t\geq 2$ and $n$ is sufficiently large, then the following holds:
@@ 5 Concluding remarks
Let $G=kP_{4}\cup tP_{2}$ for three non-negative integers $k$, $t$, and $n$, where $n\geq 4k+2t$. We define $E^{\prime}$ as a subset of the edge set $E(G)$ such that each endpoint of these edges has a degree of two in $G$. In this study, we explored the anti-Ramsey number in the context of edge deletion, demonstrating that both decreasing and unchanging behaviors are possible outcomes. We prove that the behavior of the anti-Ramsey number remains consistent when the edges in $E^{\prime}$ are removed from $G$, i.e., $AR(n,G)=AR(n,G-E^{\prime})$, if one of the conditions: (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; and (iii) $k=1$, $t\geq 1$, and $n\geq 2t+4$ holds. However, this consistency does not hold when $k\geq 2$, $t=0$, and $n=4k$. As a result, we calculate $AR(kP_{4}\cup tP_{2})$ for the cases: (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; (iii) $k=1$, $t\geq 0$, and $n\geq 2t+4$; and (iv) $k\geq 1$, $t=0$, and $n=4k$. For future research, the following questions may be considered:
@@
Let $G$ be a graph with an order of at most $n$.
@@ Acknowledgments
The research of Ali Ghalavand and Xueliang Li was supported the by NSFC No. 12131013.
Conflicts of interest
The authors declare no conflict of interest.
Data availability
No data was used in this investigation.

### v3

@@ Abstract
According to a study by Erdős et al. in 1975, the anti-Ramsey number of a graph $G$, denoted as $AR(n,G)$, is defined as the maximum number of colors that can be used in an edge-coloring of the complete graph $K_{n}$ without creating a rainbow copy of $G$. In this paper, we investigate the anti-Ramsey number under edge deletion and demonstrate that both decreasing and unchanging are possible outcomes. For three non-negative integers $k$, $t$, and $n$, let $G=kP_{4}\cup tP_{2}$. Let $E^{\prime}$ be a subset of the edge set $E(G)$ such that every endpoint of these edges has a degree of two in $G$. We prove that if one of the conditions (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; (iii) $k=1$, $t\geq 1$, and $n\geq 2t+4$, occurs then the behavior of the anti-Ramsey number remains consistent when the edges in $E^{\prime}$ are removed from $G$, i.e., $AR(n,G)=AR(n,G-E^{\prime})$. However, this is not the case when $k\geq 2$, $t=0$, and $n=4k$. As a result, we calculate $AR(n,kP_{4}\cup tP_{2})$ for the cases: (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; (iii) $k=1$, $t\geq 0$, and $n\geq 2t+4$; (iv) $k\geq 1$, $t=0$, and $n=4k$.
@@ 1 Introduction
For a positive integer $n$ and a graph $H$, the anti-Ramsey number $AR(n,H)$ is defined as the maximum number of colors in an edge-coloring of the complete graph $K_{n}$ that does not contain a rainbow copy of $H$. This concept was introduced by Erdős et al. in 1975 [5], where they established its connection to Turán numbers and conjectured values for the anti-Ramsey numbers of paths and cycles.
Subsequent research has extensively investigated anti-Ramsey numbers for edge-disjoint subgraphs, as well as for graphs composed of multiple small, disjoint components. Results concerning the anti-Ramsey numbers for spanning trees in complete graphs and more general settings can be found in [1, 3, 14, 20, 21]. The anti-Ramsey numbers for matchings (collections of disjoint edges) in complete graphs are discussed in [4, 7, 12, 14, 26], while extensions to uniform hypergraphs appear in [10, 11, 25]. More recently, research has shifted toward more intricate structures, particularly vertex-disjoint unions of small connected graphs and matchings. Our work advances this line of inquiry by examining linear forests composed of short odd paths. For clarity, we summarize the existing results most relevant to our study below. Simonovits and Sós [24] proved that if $t\geq 2$ and $n$ is sufficiently large, then the following holds:
@@ 5 Concluding remarks
Let $G=kP_{4}\cup tP_{2}$ for three non-negative integers $k$, $t$, and $n$, where $n\geq 4k+2t$. We define $E^{\prime}$ as a subset of the edge set $E(G)$ such that each endpoint of these edges has a degree of two in $G$. In this study, we explored the anti-Ramsey number in the context of edge deletion, demonstrating that both decreasing and unchanging behaviors are possible outcomes. We prove that the behavior of the anti-Ramsey number remains consistent when the edges in $E^{\prime}$ are removed from $G$, i.e., $AR(n,G)=AR(n,G-E^{\prime})$, if one of the conditions: (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; and (iii) $k=1$, $t\geq 1$, and $n\geq 2t+4$ holds. However, this consistency does not hold when $k\geq 2$, $t=0$, and $n=4k$. As a result, we calculate $AR(n,kP_{4}\cup tP_{2})$ for the cases: (i) $t\geq k+1\geq 2$ and $n\geq 8k+2t-4$; (ii) $k,t\geq 1$ and $n=4k+2t$; (iii) $k=1$, $t\geq 0$, and $n\geq 2t+4$; and (iv) $k\geq 1$, $t=0$, and $n=4k$. For future research, the following questions may be considered:
@@
Let $G$ be a graph with an order of at least $n$.
@@ Acknowledgments
The research of Ali Ghalavand and Xueliang Li was supported the by the Natural Science Foundation of China under Grant No.12131013. Jin was supported by the Natural Science Foundation of China under Grant No.12571380.